# Identifying metric structures of deep latent variable models

### v1

@@ 1 Introduction
In this paper, we change the identifiability question to arrive at working tools that do not require data labels and only impose minimal restrictions on the used models. Instead of identifying the latent variables, we identify the relationship between latent variables. For example, instead of identifying the coordinates of the latent variables we identify pairwise distances. In our experience, this matches the needs of domain experts who rarely assign meaning to the coordinates of latent variables. Using differential geometry, we prove strong identifiability guarantees on pairwise distances, angles, and more. We empirically validate our theory on two variational autoencoders.
@@ 2 Background and notation
Deep latent variable models learn densities of data $X\in\mathcal{D}$ parametrized by latent variables $Z\in\mathcal{Z}$, such that $p(X)=\int p(X|Z)p(Z)\mathrm{d}Z$ (Tomczak, 2024). We consider models with continuous latent variables, i.e., $\mathcal{Z}\subseteq\mathbb{R}^{n}$. Examples of this model class includes probabilistic PCA (Tipping & Bishop, 1999), variational autoencoders (Kingma & Welling, 2013; Rezende et al., 2014), normalizing flows (Tabak & Vanden-Eijnden, 2010; Lipman et al., 2022), diffusion models (Ho et al., 2020) and more.
@@ 3.1 Assumptions on $\mathcal{F}$ and $\mathcal{Z}$
    The differential of $f$, $\mathrm{d}f$, has full column rank.
@@
    All $f\in\mathcal{F}$ have the same image. That is, for any $f_{a},f_{b}\in\mathcal{F}$, we have $f_{a}(\mathcal{Z})=f_{b}(\mathcal{Z}):=\mathcal{M}\subseteq\mathcal{D}$.
@@ 5 Experiments
Theorem 4.7 states that geodesic distances are identifiable while Euclidean ones are not. This suggests that geodesic distances should be more stable under model retraining than the Euclidean counterpart. To test this hypothesis, we train 30 models with different initial seeds and compute both Euclidean and geodesic distances between 100 randomly chosen point pairs from the test set. We emphasize that the pairs are the same across all models, which allows us to measure the variances of the distances across models.
@@
To assess the stability of the distance measures, we compute the coefficient of variation for each pair, which is evaluated as $\text{CV}(m_{pair})=\nicefrac{{\sigma(m_{pair})}}{{\mu(m_{pair})}}$ with $\sigma$ and $\mu$ denoting standard deviation and the mean, respectively. The coefficient of variation is a unitless measure of variability, where low values indicate less variability. We use this to compare the variability of Euclidean and geodesic distances.
We consider two models and datasets. First, a smaller model that satisfies all our assumptions, and second a larger model where we disregard the injectivity assumption. To compute geodesics we parametrize them by a spline connecting two points in the latent space and minimize its energy. The discussion around Definition A.23 in Appendix A covers how this leads to a geodesic. It has been noted that taking decoder uncertainty into account is key to good performance (Arvanitidis et al., 2018; Hauberg, 2018) and we follow the ensemble-based approach from Syrota et al. (2024), implying that we train an ensemble of 8 decoders. Details are in Appendix C and the code to reproduce our results is available at https://github.com/mustass/identifiable-latent-metric-space/.
mnist with A1-A4 satisfied.   We use $\mathcal{M}$-flows (Brehmer & Cranmer, 2020) to construct a VAE with an injective decoder. We train this model on a 3-class subset of mnist with a 2D latent space for visualization purposes. An example of a geodesic curve from digit 7 to digit 0 from the test set is visualized in Fig. 5. The geodesic crosses class boundaries where they are well-explored by the model and offer little uncertainty.
The left panel of Fig. 7 shows a histogram of the coefficient of variation for the 100 point-pairs, where we see a narrower distribution with both a lower mean and spread for geodesic distances. We perform a one-sided Student’s $t$-test for the null hypothesis that geodesic distances vary less than the Euclidean (Table 1) and find strong evidence for the hypothesis. This demonstrates that identifiability improves reliability.
@@
celeba with A2 relaxed and A3 verified.   The general VAE model is known to be effective due to its flexible decoder parametrized by a neural network with arbitrary architecture that is not guaranteed to be globally injective (A2) nor to have full rank Jacobian (A3). We train this model on the celeba dataset (Liu et al., 2015) where the architecture is composed of convolutional and dense layers with ELU activation functions. The latent space dimension is 64 and we further employ Resize-Conv layers (Odena et al., 2016) to improve image quality. We follow the approach from Shao et al. (2018) to validate that the decoder Jacobian is, indeed, always full rank. An example geodesic is shown in Fig. 6 alongside a Euclidean counterpart, where we do not observe a significant difference between generated images.
Fig. 7(right) shows that the coefficient of variation for geodesic distances has both lower mean and standard deviation than Euclidean distances. The one-sided Student’s $t$-test again validates this observation (Table 1). This demonstrates that geodesic distances remain more reliable than Euclidean ones even when the injectivity assumption may be violated.
mnist celeba
$t$-statistic -8.64 -22.33
$p$-value 1.00 1.00
@@
Figure 7: Histograms of coefficients of variation for Euclidean and geodesic distances on mnist (left) and celeba (right). Geodesic distances vary significantly less, which is quantified in Table 1.
@@ 6 Relations to existing work
Relative representations consider pairwise latent Euclidean distances as a form of representation (Moschella et al., 2022). This is well-aligned with the ideas we put forward. Empirically, Moschella et al. (2022) reports that latent Euclidean distances are sufficiently robust to be of practical value. We find that identifiable geodesic distances are significantly more robust, such that our approach should bring both mathematical guarantees and empirical improvements.
@@ Acknowledgments
JX is supported by an Natural Sciences and Engineering Research Council of Canada (NSERC) Canada Graduate Scholarship. BBR acknowledges the support of NSERC: RGPIN2020-04995, RGPAS-2020-00095.
@@ C.2 Optimizing the spline to find a geodesic
where $\Delta_{t}=\bar{t}_{i}-\bar{t}_{i-1}$ is the step size in the discretization of the interval $[0,1]$ and is assumed to be constant. The decoders $f_{\hat{\theta}_{k}}$ and $f_{\hat{\theta}_{kj}}$ are sampled uniformely and independently from the ensemble.

### v3

@@ 1 Introduction
In this paper, we change the identifiability question to arrive at working tools that do not require data labels and only impose minimal restrictions on the used models. Instead of identifying the latent variables, we identify the relationship between latent variables. For example, instead of identifying the coordinates of the latent variables we identify pairwise distances. In our experience, this matches the needs of domain experts who rarely assign meaning to the coordinates of latent variables. Using differential geometry, we prove strong identifiability guarantees on pairwise distances, angles, and more. We empirically validate our theory on four different generative models.
@@ 2 Background and notation
Deep latent variable models learn densities of data $X\in\mathcal{D}$ parametrized by latent variables $Z\in\mathcal{Z}$, such that $p(X)=\int p(X|Z)p(Z)\mathrm{d}Z$ (Tomczak, 2024). We consider models with continuous latent variables, i.e., $\mathcal{Z}\subseteq\mathbb{R}^{n}$. Examples of this model class include probabilistic PCA (Tipping & Bishop, 1999), variational autoencoders (Kingma & Welling, 2013; Rezende et al., 2014), normalizing flows (Tabak & Vanden-Eijnden, 2010; Lipman et al., 2022), diffusion models (Ho et al., 2020) and more.
@@ 3.1 Assumptions on $\mathcal{F}$ and $\mathcal{Z}$
    The differential of $f$, $\mathrm{d}f$, has full column rank
@@
    All $f\in\mathcal{F}$ have the same image. That is, for any $f_{a},f_{b}\in\mathcal{F}$, we have $f_{a}(\mathcal{Z})=f_{b}(\mathcal{Z}):=\mathcal{M}\subseteq\mathcal{D}$
@@ 5 Experiments
Identifiability is a theoretical concept studied in the asymptotic regime of infinite data. Our experiments aim to demonstrate that this asymptotic property is practically exploitable in standard models using off-the-shelf methods.
Theorem 4.7 proves that geodesic distances are identifiable, while Euclidean ones are not. This suggests that geodesic distances should be more stable under model retraining than the Euclidean counterpart. To test this hypothesis, we train 30 models with different initial seeds and compute both Euclidean ($d_{E}$) and geodesic ($d_{g}$) distances in the latent space between 100 randomly chosen unique point pairs ($\left\{p_{i}=(\mathbf{x}_{j},\mathbf{x}_{k})\right\}_{i=1}^{100},\mathbf{x}_{j},\mathbf{x}_{k}\in\mathcal{M},\text{ }j\neq k$) from the test set. We emphasize that the pairs are the same across all models, allowing us to measure the variances of the distances across models.
@@
To assess the stability of the distance measures, we compute the coefficient of variation for each pair, which is evaluated as $\text{CV}(d_{*}(p_{i})))=\nicefrac{{\sigma(d_{*}(p_{i}))}}{{\mu(d_{*}(p_{i})))}}$ with $\sigma$ and $\mu$ denoting standard deviation and the mean across 30 seeds, and $d_{*}$ is a placeholder for the distance measure used ($d_{E}$ or $d_{g}$). The coefficient of variation is a unitless measure of variability, where low values indicate less variability. We use this to compare the variability of Euclidean and geodesic distances.
We consider two models and four datasets. First, a model that satisfies all our assumptions, and second, a model where we disregard the injectivity assumption. To compute geodesics we parametrize them by a spline connecting two points in the latent space and minimize its energy. The discussion around Definition A.23 in Appendix A covers how this leads to a geodesic. It has been noted that taking decoder uncertainty into account is key to good performance (Arvanitidis et al., 2018; Hauberg, 2018) and we follow the ensemble-based approach from Syrota et al. (2024), implying that we train an ensemble of 8 decoders. Details on computing geodesics and experimental details are in Appendices C and D. The code to reproduce our results is available in the project repository GitHub [^1] https://github.com/mustass/identifiable-latent-metric-space.
mnist and cifar10 with A1-A4 satisfied.   We use $\mathcal{M}$-flows (Brehmer & Cranmer, 2020) to construct a model with an injective decoder. We train this model on a 3-class subset of mnist (Deng, 2012) with a 2D latent space for visualization purposes and full cifar10 (Krizhevsky et al., ). An example of a geodesic curve from digit 7 to digit 0 from the test set is visualized in Fig. 5. The geodesic crosses class boundaries where they are well-explored by the model and offer little uncertainty.
The left side of Fig. 7 shows a histograms of the coefficient of variation for the 100 point-pairs, where we see a narrower distribution with both a lower mean and spread for geodesic distances. We perform a one-sided Student’s $t$-test for the null hypothesis that geodesic distances vary less than the Euclidean (Table 1) and find strong evidence for the hypothesis. This demonstrates that identifiability improves reliability.
@@
fmnist and celeba with A2 relaxed and A3 verified.   The general VAE model is known to be effective due to its flexible decoder parametrized by a neural network with arbitrary architecture that is not guaranteed to be globally injective (A2) nor to have full rank Jacobian (A3). We train this model on the fmnist (Xiao et al., 2017) and celeba (Liu et al., 2015) datasets where the architecture is composed of convolutional and dense layers with ELU activation functions. The latent space dimension is 64 and we further employ Resize-Conv layers (Odena et al., 2016) to improve image quality. We follow the approach from Shao et al. (2018) to validate that the decoder Jacobian is, indeed, always full rank. An example geodesic is shown in Fig. 6 alongside a Euclidean counterpart, where we do not observe a significant difference between generated images.
Figs. 7(b) and 7(d) (right side) show that the coefficient of variation for geodesic distances has both lower mean and standard deviation than Euclidean distances. The one-sided Student’s $t$-test again validates this observation (Table 1). This demonstrates that geodesic distances remain more reliable than Euclidean ones even when the injectivity assumption may be violated.
mnist fmnist cifar10 celeba
$t$-statistic -8.64 -16.75 -42.83 -22.33
$p$-value 1.00 1.00 1.00 1.00
@@
(a) $\mathcal{M}$-flow model on mnist
(b) VAE on fmnist
(c) $\mathcal{M}$-flow model on cifar10
(d) VAE on celeba
Figure 7: Histograms of coefficients of variation for Euclidean and geodesic distances on mnist (7(a)), fmnist (7(b)), cifar10 (7(c)) and celeba (7(d)). Geodesic distances vary significantly less, which is quantified in Table 1.
@@ 6 Relations to existing work
Transformations of the latent space have proven valuable in various areas, including disentangled and equivariant learning, highlighting their mathematical and conceptual connections. Higgins et al. (2018); Wang et al. (2021); Zhu et al. (2021) assume a disentangled latent space where transformations decompose into individual factors of variation and the goal is to find representations that respect this structure. This can be interpreted as equivariance of the generator transformations $A_{a,b}$ (Definition 4.2). Instead of enforcing specific properties of $A_{a,b}$, our theory analyzes the natural properties of the generator transformations, and we find that just by learning a generative model, $A_{a,b}$, will automatically respect the latent Riemannian geometry (Theorem 4.5). Thus, our theory is valid regardless of whether a disentangled latent space exists in the sense of Higgins et al. (2018).
@@
Relative representations are explicitly constructed from the original representations and a set of anchor points to be invariant under angle-preserving transformations (using cosine similarity) (Moschella et al., 2022) or isometries (using a proxy for geodesic distances) Yu et al. (2024). Our approach instead provides mathematical guarantees on invariance under isometries of the original representations and establishes the link to statistical identifiability.
@@ Acknowledgments
This work was supported by a research grant (42062) from VILLUM FONDEN. This project received funding from the European Research Council (ERC) under the European Union’s Horizon research and innovation programme (grant agreement 101125993). The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). JX is supported by an Natural Sciences and Engineering Research Council of Canada (NSERC) Canada Graduate Scholarship. BBR acknowledges the support of NSERC: RGPIN2020-04995, RGPAS-2020-00095.
@@ C.2 Optimizing the spline to find a geodesic
where $\Delta_{t}=\bar{t}_{i}-\bar{t}_{i-1}$ is the step size in the discretization of the interval $[0,1]$ and is assumed to be constant. The decoders $f_{\hat{\theta}_{k}}$ and $f_{\hat{\theta}_{kj}}$ are sampled uniformly and independently from the ensemble.
Appendix D Experiments
$\mathcal{M}$-flow on mnist and cifar10
The decoder is modeled by 10 coupling flow layers introduced by (Dinh et al., 2015) where the conditioner is 2-layer MLP with 50 hidden units and ReLu activation function, and transformer is the RQS spline (Durkan et al., 2019) in the interval $[-3,3]$ with 8 knots (22 knots for cifar10). The decoder has 5.5 million parameter for mnist and 55 million for the cifar10 model.
The encoder consists of 4 Residual Blocks (He et al., 2015) with Convolutional transformations each time multiplying the number of channels by 5 and the Relu activation function. The last layer is the linear layer with hyperbolic tangent function multiplied by 3 at the end to constrain the output to [-3,3] and accommodate the coupling flow above. The encoder has 135 thousand parameters for both mnist and cifar10 models.
Adam (Kingma & Ba, 2017) was used for all the experiments.
VAE on fmnist and celeba
The encoder consists of five $4\times 4$ convolutional layers, the first two with 128 channels, followed by the next three with 256 channels, with the first and last having a stride of one and the others a stride of two; two linear layers, one with 256 units and the second with 64 units (latent dimensions) for the VAE approximate posterior mean and 64 for the VAE approximate posterior variance.
The decoder starts with two fully-connected layers, one with 256 units and one with 16384 units; five $4\times 4$ resize convolutional layers (Odena et al., 2016), with the same stride configuration as the encoder, but with 256, 128, 64, and 32 channels, respectively. All layers in the encoder and decoder have an ELU activation function. Total number of parameters: $6.5$ million.
The VAE is trained with an additional loss term coming from the Perceptual loss (Johnson et al., 2016).
Geodesic computation
The geodesic computation was done on models that had an ensemble of 8 decoders following the procedure of (Syrota et al., 2024) and hyperparameters in Table 2.
Parameter Value
Initialization of free parameters zeros (straight line)
Number of polynomials in the spline 10
Discretization in time (energy) 256
Discretization in time (final length) 256
Optimizer Adam (Kingma & Ba, 2017)
Max. steps 4096
Learning rate 0.01
Early stopping patience (steps) 100
Early stopping delta 1.0
Table 2: Shared geodesics training parameters for all models.